Boating: A family travels 18 miles down river and returns. It takes 8 hours to make the round trip. Their rate in still water is twice the rate of the rivers current. How long will the return trip take?
Answer by psbhowmick(878)

Let the river current be x km/hr.
Then, speed of the boat is 2x km/hr.

Time taken to go 18km downstream = 18/(2x+x) = 6/x hr.

Time taken to go 18km upstream = 18/(2x-x) = 18/x hr.

Total time taken = 6/x + 18/x hr.
6/x + 18/x = 8

24/x = 8

x = 3

Therefore, time taken for return trip (18km upstream) = 18/(2x-x) = 18/3 = 6 hr.